OpenStudy (anonymous):

Prove that if f(x) = integral from 0 to x of f(t) dt then f = 0

OpenStudy (anonymous):

if \[f(x) = \int\limits_{0}^{x} f(t) dt \] then f = 0

OpenStudy (mr.math):

This is not a true statement.

OpenStudy (anonymous):

Michael Spivak claims it is

OpenStudy (mr.math):

Who's Michael Spivak?

OpenStudy (anonymous):

The man who wrote my textbook.

OpenStudy (mr.math):

Whoever he might be, tell him Newton has another opinion :P

OpenStudy (mr.math):

This is a direct use of the fundamental theorem of calculus.

OpenStudy (anonymous):

\[\int_0^x 2t dt= t^2 \] \[x^2\]

OpenStudy (mr.math):

Oh wait!

OpenStudy (anonymous):

haha im not going anywhere with this one

OpenStudy (across):

It implies f(0)=0.

OpenStudy (anonymous):

hes claiming f(x) = 0 for any x

OpenStudy (anonymous):

I certinally dont see it

OpenStudy (anonymous):

is there derivative sign infront of integral?

OpenStudy (anonymous):

Nope that is the whole question

OpenStudy (mr.math):

I didn't read the question well at first. This means that f is an anti-derivative of itself, if I'm seeing this right.

OpenStudy (anonymous):

yes f'(x) = f(x)

OpenStudy (across):

We know that\[f(x)=\int f(x)dx\implies f(x)=e^x\]^^

OpenStudy (mr.math):

Yeah.

OpenStudy (anonymous):

oh nice

OpenStudy (anonymous):

I didnt think about e^x with this one

OpenStudy (mr.math):

Then differentiate both sides you get, f'(x)=f(x), which is an ODE that has the solution \(f(x)=ce^{x}\).

OpenStudy (mr.math):

Your statement is still not correct :P

OpenStudy (anonymous):

Welllll hold on

OpenStudy (zarkon):

it is correct....find c

OpenStudy (anonymous):

the e^x makes senese with f'(x) = f(x) but this function is an integral

OpenStudy (mr.math):

You're saying c=0 @Zarkon.

OpenStudy (zarkon):

\[f(0) = \int\limits_{0}^{0} f(t) dt=0\]

OpenStudy (anonymous):

It wants me to prove the function is 0

OpenStudy (zarkon):

\[\Rightarrow c=0\]

OpenStudy (mr.math):

That's right! I'm a loser!! :(

OpenStudy (anonymous):

No you're not! So basically because f'(x) = f(x) I can say f(x) = ce^x and then show that f(0) = 0 implying that c=0 so f = 0

OpenStudy (mr.math):

Exactly!

OpenStudy (anonymous):

makes perfect sense! Im just be a loser now but how do we know there is no other function s.t. f'(x) = f(x)

OpenStudy (across):

Mr. Michale Spivak did a good job, that is, to elicit eager students to congregate and think this one through. xd

OpenStudy (anonymous):

I hate michale spivak :P

OpenStudy (anonymous):

I'm just kidding its just a challening course for me

OpenStudy (mr.math):

Lol @across. @Wall, this is a first order homogeneous equation and had only this solution, \(i.e f(x)=ce^{x}\).

OpenStudy (mr.math):

homogeneous differential equation* and it has* *_*

OpenStudy (anonymous):

ahhh yes! I really like this problem! I can't believe I didnt realize f(x) had to be some form of e^x Thanks for all the help everyone!!!!

OpenStudy (mr.math):

You're welcome! Thanks for fanning me :D